How many atoms are in
37.9
37.9
g of potassium bromite (KBrO2)
1
Expert's answer
2019-07-19T03:25:24-0400

find the molar mass of the substance: M(KBrO2)=M(K)+M(Br)+2M(O)=151g/mol

Now find the number of mollecules of the substance: Nm=(m(KBrO2)/M(KBrO2))*NA1.5*1023 (NA -Avogadro constant )

In this mollecule, the number of atoms is 4 times greater than the number of mollecules: N=Nm*4=6*1023;

Answer: 6*1023